There is a Circle with the Diameter as 5. This is inside of a square which has a base of 17. How Do You Find the Area? Please show all work
Answer by nyc_function(2741)

Find the area of what? The square? The circle? The shaded region? If you seek the area of the shaded region, which region?
If you simply want the area of the circle, then we need to know certain facts from the given information.
If the diameter is 5, then the radius is (1/2) times diameter or in this case
(5/2).
We now use the area of a circle formula, which is A = pi*(r^2), where pi = 3.14
and r = (5/2).
A = (3.14)(5/2)^2
A = 19.625
Is this what you are looking for?